Write the series with summation notation: 2+5+8+11+...
Found 2 solutions by stanbon, jim_thompson5910:
Answer by stanbon(75887)

Write the series with summation notation: 2+5+8+11+...
2
2+1*3 = 5
2+2*3 = 8
2+3*3 = 11
---
a(n) = 2 + 3n
---
Answer: S(n) = Summation (as n goes from 0 to k) of (2+3n)
-------------
Cheers,
Stan H.
-------------

Answer by jim_thompson5910(35256)

Method #1

-----------------------------------------------------------------------------
Method #2

Notice how we add 3 to each term to get the next term (2+3 = 5, 5+3 = 8, etc). This points to the fact that we have an arithmetic sequence.

So the common difference is d = 3

The first term is a(1) = 2

Use these two facts to get the nth term

a(n) = a(1) + d(n-1)
a(n) = 2 + 3(n-1)
a(n) = 2 + 3n-3
a(n) = 3n - 1

So because the nth term is 3n-1 this means that 2+5+8+11+... = \sum_{k=1}^{\infty}(3k-1) (the only difference is that I'm using k instead of n as the index)

------------------------------------------------------------------------------------------------------------------------

If you need more one-on-one help, email me at jim_thompson5910@hotmail.com. You can ask me a few more questions for free, but afterwards, I would charge you ($2 a problem to have steps shown or $1 a problem for answer only).

Alternatively, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Any amount is greatly appreciated as it helps me a lot. This donation is to support free tutoring. Thank you.

Jim
------------------------------------------------------------------------------------------------------------------------